Write an equation of the line that is parallel to the graph of y-3x=4 and has the same y- intercept as the graph for 4y+x=36
Answer by algebrahouse.com(1659)

"Write an equation of the line that is parallel to the graph of y-3x=4 and has the same y- intercept as the graph for 4y+x=36"

*Notes*
- Parallel lines have equal slopes.
- Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

y - 3x = 4 {need to be parallel to this line}
y = 3x + 4 {changed to slope-intercept form}
slope = 3

4y + x = 36 {need to have same y-intercept as this line}
4y = -x + 36 {subtracted x from both sides}
y = -1/4x + 9 {changed to slope-intercept form}
y-intercept = 9

For the new line, the slope is 3 and the y-intercept is 9
y = 3x + 9 {substituted slope and y-intercept into slope-intercept form}
www.algebrahouse.com